• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name jroche

    Log entry time 11:39:00 on September 15, 2010

    Entry number 332265

    Followups:

    keyword=LHRS detector map testing :status

    At this point, here is the status of the detectors in the left hrs: -One of the S2 right tube is giving no signal. The HV interface shows that this tube draws no current. Hisham and I will try first to plug unplug the HV wire from the PMT base. Note the HV supply map for these detectors do not match the scaler map. Do we care?

    -I have not seen any meaningful signal from the S2R ADCS since I started looking at cosmics. First the cables were not plugged, Alex and I plug them earlier this week. From the xscaler display and the TDC I know that there are signal there. For these tubes I measure a pedestal that is wider than the one of their counter parts on the left. Roughly a factor of 30 times wider. I tried to invert the signal polarity in the ADC yesterday: it did not work. Back to the drawing board.

    -Yesterday Hisham plugged the TDC signals for S2, S1 and the cerenkov in high resolution TDCS. I updated the detector map and ran the code. Apparently I need to do more modifications because I am getting no TDC info for these detectors. I'll investigate that today.

    -Cerenkov: right now triggering on cosmics the cerenkov are reading out 2kHz. That's sound pretty high to me. Does that makes sense?



    A copy of this log entry has been emailed to: camsonne@jlab.org, munoz@jlab.org, hyde@jlab.org